OpenStudy (anonymous):

The rectangles shown are similar. What is the value of x? A. 7 B. 8 C. 6 D. 12

OpenStudy (anonymous):



OpenStudy (anonymous):

I think D.

OpenStudy (radar):

Think again.

OpenStudy (igreen):

We set up a proportion. \(\dfrac{x}{24} = \dfrac{5}{15}\) Multiply 24 to both sides: \(\dfrac{24 \times 5}{15}\)

OpenStudy (anonymous):

Or... 6!

OpenStudy (igreen):

Simplify that and you get your answer..

OpenStudy (anonymous):

8?

OpenStudy (radar):

Now you're thinking

OpenStudy (igreen):

Yep, 8 is correct.